Circular No. 7159

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999by IN NGC 2841 P. Garnavich, S. Jha, and R. Kirshner, Harvard-Smithsonian Center for

Astrophysics, write: “Spectra obtained by P. Berlind with the 1.5-m Tillinghast telescope on May 6.2 UT confirm that SN 1999by is a type-Ia event, but the spectrum is peculiar. The 580-nm Si II feature is very strong relative to the 615-nm Si II absorption, suggesting that this is a subluminous type-Ia supernova. The ratio of the depth of 580- to 615-nm lines correlates with the decline rate (Nugent et al. 1995, Ap.J. 455, 147) and is found to be 0.6 in SN 1999by. This is similar to the ratio observed in SN 1991bg, the prototypical fast-declining, subluminous type-Ia supernova. In addition, the flux around 400 nm is depressed relative to normal type-Ia events, consistent with Ti II absorption also seen in subluminous events. The minimum of the Si II absorption, corrected for the redshift of NGC 2841 (638 km/s), indicates a photospheric velocity of 10 800 km/s. Spectra of SN 1999by can be found

at http://cfa-www.harvard.edu/cfa/oir/Research/supernova.html.”

V364 PEGASI Q. Y. Qiao, J. Y. Wei, Y. L. Qiu, and J. Y. Hu, Beijing Astronomical

Observatory (BAO), report their discovery in the course of the BAO supernova survey of an outburst of the cataclysmic variable in Pegasus that was previously reported on IAUC 6746, 6772, 6982, and 7021. The unfiltered CCD magnitude of the star was 15.0 on May 5.82 UT.

V4334 SAGITTARII T. Geballe, Gemini North, Joint Astronomy Centre; S. Eyres, Liver-

pool John Moores University; and A. Evans and B. Smalley, Keele University, report on infrared (JHKLM) spectra obtained on Apr. 22 and May 3 UT: “There is now strong He I emission at 1.083 μm, in contrast to the strong absorption in this line in 1998 (cf. IAUC 6896; Eyres et al., MNRAS, in press). No other emission or absorption lines are obviously present. There is also a strong dust continuum, at temperature ∼ 1100 K (cf. IAUC 7155). Longward of 1 μm, the dust emission now completely obscures the stellar photosphere, which was clearly visible in 1998.”

CORRIGENDUM On the last two lines of IAUC 7158, the Apr. 13.675 and 16.662 obser-

vations attributed to SN 1999be actually belong to SN 1999bg in IC 758.

1999 May 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green